What is the molarity of a solution in which 23 grams of NaCl is dissolved in 1.08 L solution , with work shown.
1
Expert's answer
2020-04-10T13:35:26-0400

Solution:

By definition, molarity = moles of solute / volume of solution.

Mr(NaCl) = 58.443 g/mol. Then,

1 mol of NaCl has a mass of 58.443 g, so

Moles of NaCl = n(NaCl) = (23 g NaCl) * (1 mol NaCl / 58.443 g NaCl) = (23 / 58.443) = 0.3935 moles.

Volume of solution = V(solution) = 1.08 L.

Molarity of NaCl = C(NaCl) = n(NaCl) / V(solution) = (0.3935 moles) / (1.08 L) = 0.3644 M.

Molarity of NaCl = 0.3644 M.


Answer: Molarity of a solution of NaCl is 0.3644 M.